OpenStudy (fanduekisses):

[PRECAL HELP~*] Write the sum using summation notation, assuming the suggested pattern continues. -9 - 4 + 1 + 6 + ... + 66

OpenStudy (mathmale):

Is this an arithmetic sequence or a geometric sequence? How do you know? If you start with -9 and find that your next term in the sequence is -4, what had you done to that first term (-9)?

OpenStudy (fanduekisses):

added positive 5

OpenStudy (mathmale):

that's right! So, have we an arithmetic or a geom. sequence here?

OpenStudy (fanduekisses):

arithmetic

OpenStudy (mathmale):

Very good! Now, please look up "sum of an arithmetic sequence" anywhere you think appropriate: in your textbook, in your online learning materials, through an Internet search, whatever. by doing this you'll find a formula for the sum of an arithmetic sequence of n elements.

OpenStudy (fanduekisses):

is it this one?

OpenStudy (fanduekisses):

or this one

OpenStudy (mathmale):

Yes and no. If you were to find all of the terms of the sequence and then add them together, you'd get the sum of the sequence. But there's another formula that requires you to add only the first and last terms of the sequence and to multiply the result by n/2, where n is the number of terms you're adding together. What is that formula? Hint: see: http://www.purplemath.com/modules/series4.htm

OpenStudy (fanduekisses):

so -9 + 66 (5/2) ?

OpenStudy (mathmale):

First, are you sure that there are only 5 terms in this sequence? Secondly, it's very important that you enclose that -9+66 inside parentheses. Otherwise others would think you're multiplying only 66 by (5/2).

OpenStudy (fanduekisses):

ohh I forgot about the (...)

OpenStudy (mathmale):

I see you've counted the number of visible terms and decided that n = 5 based upon that. But that's not right. Hint: There's another very useful formula that applies here: l=a + (n-1)d. Have you seen this before?

OpenStudy (fanduekisses):

no

OpenStudy (mathmale):

a= first term l= last term n= number of terms (which is what you want) d= the number you keep on adding to each most recent term to get the next term.

OpenStudy (mathmale):

At this point, you have a choice: 1) write out all the terms, from -9 through 66, or 2) use this formula I've given you, to find a) n-1 and b) n.

OpenStudy (fanduekisses):

66= -9 + (n-1)

OpenStudy (fanduekisses):

75=n-1

OpenStudy (mathmale):

Hint: don't forget that ' d '. ' d ' = 5,

OpenStudy (fanduekisses):

75= (n-1) 5 15=(n-1)

OpenStudy (mathmale):

Nice work! So, n = ??

OpenStudy (fanduekisses):

16?

OpenStudy (mathmale):

Great! Your arith. sequence has 16 terms. Now go back to the formula for calculating the sum of the first n terms: (a + l)(n/2).

OpenStudy (mathmale):

a=-9 d=? l=? n=?

OpenStudy (fanduekisses):

(-9 +66) (16/2) = 456?

OpenStudy (mathmale):

Surely looks good! Congrats. Any questions?

OpenStudy (fanduekisses):

Thank you :D though the answer options are a little different:

OpenStudy (mathmale):

In those answer options, you see the Greek letter sigma in each: \[\sum_{}^{}\] what does that represent in this arithmetic sequence problem?

OpenStudy (fanduekisses):

the sum

OpenStudy (mathmale):

Right. That "sigma" represents "summation." Now look at the lower limit and the upper limit on each sigma. Supposing n (a counter, like counting on your fingers) starts at 0 and ends at 15. How many terms would you be adding up?

OpenStudy (fanduekisses):

16

OpenStudy (mathmale):

Right. See what I'm driving at? Does this give you enuf info to choose the correct answer, or would you like to discuss further how to identify the correct answer?

OpenStudy (fanduekisses):

is it the second one?

OpenStudy (mathmale):

Yes. You could check your answer as follows: -9+0(5) =-9 -9+1(5) =-4 -9+2(5) = 1 and so on.... and you'd find that the sum of these 16 terms is 456. Again, nice work!

OpenStudy (mathmale):

I need to get off the 'Net now, but look forward to working iwth you again in the future!

OpenStudy (fanduekisses):

Thanks you very much for your wonderful help <3